# No Validation, No Problem: Predicting Model Performance from a Single Gradient

### v1

@@ Abstract
We propose a validation‑free checkpointing signal from a single forward–backward pass: the Frobenius norm of the classifier‑head gradient on one detached‑feature batch, $\|g\|_{F}=\|\nabla_{W}\mathcal{L}\|_{F}$. Across ImageNet‑1k CNNs and Transformers, this proxy is strongly negative with Top‑1 and positive with loss. Selecting the checkpoint with the minimum head gradient in a short tail window closes most of the gap to the oracle ($4.24\%\!\pm\!2.00\%$ with a universal setup, $\approx\!1.12\%$ with light per‑family tuning). For practical deployment, a head‑scale normalization is more stable within classic CNN families (e.g., ResNets), while a feature‑scale normalization works well for Transformers and modern CNNs. The same one‑batch probe also predicts COCO detection/segmentation mAP. In diffusion (UNet/DDPM on CIFAR-10), it tracks progress and enables near-oracle tail-window selection; it is positively correlated with the same-distribution probe MSE and negatively with FID (lower is better), so it can be used as a lightweight, label‑free monitor. Validation labels are never used beyond reporting. The probe adds $\ll 0.1\%$ of an epoch and works as a drop‑in for validation‑free checkpoint selection and early stopping.
@@ 1 Introduction
Modern training loops heavily lean on held-out validation sets to decide when to stop and which checkpoint or model to keep. This practice costs extra compute for periodic evaluation, can be impossible when labels are scarce or private, and may be unreliable under distribution shift. We revisit a minimalistic question: can one predict model quality using just a single mini-batch and one backpropagation pass, without any validation data at selection time?
In this work, we find that the classifier head gradient carries a stable, actionable signal. Measuring the head gradient norm on a single batch (denoted as the head gradient norm $\|g\|_{F}$) yields a ranking that tracks model quality remarkably well: lower $\|g\|_{F}$ tends to indicate better Top-1 (the standard top-1 classification accuracy; i.e., the fraction of samples for which the model’s highest-probability class matches the ground truth). In practice, we run one forward/backward pass, backpropagating through the head only (features are detached), and use $\|g\|_{F}$ as a drop-in score to select checkpoints or pre-screen architectures. No validation set is required during training, but evaluation is performed on the validation set.
The probe is label-free at selection time (no validation labels), architecture-agnostic (only assumes a classifier head) and extremely cheap (a single-batch head-only backward pass). This makes it attractive in regimes where validation sets are unavailable/undesirable, and in pipelines that must evaluate many checkpoints or candidates (e.g., early stopping and lightweight NAS pre-screening).
Across three standard vision tasks, a single-batch head-gradient norm ($\|g\|_{F}$) is a reliable, training-free proxy for model quality. On ImageNet-1k (25 CNN/Transformer models), $\|g\|_{F}$ strongly anti-correlates with Top-1 ($|r|\approx-0.84$ on average; lower is better) and positively correlates with loss, with the same trend within each family and robustness to leave-one-out outliers. In COCO detection (off-the-shelf detectors), $\|g\|_{F}$ shows a clear negative association with mean Average Precision (mAP); in COCO instance segmentation (12 models from Mask2Former, Mask R-CNN and YOLO-Seg), it exhibits a strong monotonic negative correlation with mask AP (e.g., $\rho\approx-0.98$). As a practical benefit, using $\|g\|_{F}$ to pre-screen architectures in a small NAS setting (K=3 of 8) achieves high recall [32] of the best model while saving $\approx 60\%$ computation. Ablations contrasting $\ell_{1}/\ell_{\infty}/$scaled norms and Fisher-style [3] variants produce comparable averages; the plain $\ell_{2}$ head norm remains the simplest and a strong default.
Our selection rule itself never uses validation labels. To report effect sizes, we evaluated Top-1/mAP and correlations on standard validation data. We report across‑architecture correlations as macro evidence. For detection, we also report the sensitivity to evaluation size and label assignment choices. Compute and memory footprints are summarized for reproducibility and cost comparisons.
[…]
Contributions.
  • A one-shot, validation-free selection signal. We propose using the single-batch head gradient norm $\|g\|_{F}$ for checkpoint/model ranking and early stopping, requiring only one forward/backward pass on the head.
  • Robust across architectures, datasets, and tasks. Through cross-model, family-wise, and leave-one-out analyses, we show that a single-gradient statistic generalizes from ImageNet-1k [8] classification to COCO [27] detection/instance segmentation and diffusion generative modeling (UNet/DDPM [18]), it is strongly negatively associated with final accuracy/AP and positively associated with loss (or FID/probe MSE in diffusion), consistently across model families.
  • Immediate utility for practice. A simple “min-$\|g\|_{F}$” rule closely tracks the best Top-1 over training and serves as a low-cost pre-screen in NAS-style selection, and in diffusion enables label-free tail-window checkpoint selection with near-zero gap to an oracle, delivering substantial compute savings with high recall of the best candidates.
  • Low cost and easy adoption. The probe is architecture-agnostic, adds negligible overhead relative to one training epoch, and integrates into existing code with a few lines (no head refits or extra data).
@@ 2 Related Work
Connections between loss landscape geometry and generalization have been studied via flatness/sharpness and Hessian spectra (e.g., 21, 20, 39, 14). Large gradients often co-occur with sharper regions of the loss; our findings align with this picture: models/checkpoints with smaller head gradients tend to exhibit better Top-1 and lower loss, suggesting improved separability of the learned features under the current head.
@@
3 Theoretical Motivation
We build our method on two key observations:
Observation 1: Gradient Norm and Loss Curvature. The gradient norm at a checkpoint reflects the local loss curvature. For a model with parameters $\theta$ at checkpoint $t$:
@@ 3 Theoretical Motivation
Observation 2: Sharpness and Generalization. Keskar et al. [21] and Jiang et al. [20] show that flat minima (low sharpness) generalize better than sharp minima (high sharpness). This suggests:
High sharpness $\implies\text{Large gradient norm}$ (2)
$\implies\text{Poor generalization.}$ (3)
Our Insight. We posit that the classification head gradient norm, measured on training data, serves as a proxy for the model’s position in the loss landscape. Models approaching flat minima exhibit smaller gradients and better generalization, explaining the observed negative correlation (r=0.85r=-0.85).
4 Method
Setup and notation.
@@ 4 Method
Let $f_{\theta}(x)=h_{W}(\phi_{\psi}(x))$, where $\phi_{\psi}:\mathcal{X}\!\to\!\mathbb{R}^{d}$ is a feature extractor and $h_{W}(z)=Wz$ is a linear head with $W\in\mathbb{R}^{C\times d}$. (We omit the bias for clarity; if present, the bias gradient is $\nabla_{b}\mathcal{L}=\tfrac{1}{B}(P-\widetilde{Y})\mathbf{1}$ with $\mathbf{1}\!\in\!\mathbb{R}^{B}$ the all-ones vector, and we report norms on $W$ only.) Given a mini-batch $\{(x_{i},y_{i})\}_{i=1}^{B}$ of $B$ samples, define the detached feature matrix $Z=[\phi_{\psi}(x_{1}),\ldots,\phi_{\psi}(x_{B})]\in\mathbb{R}^{d\times B}$, the logits $S=WZ$, the probabilities $P=\mathrm{softmax}(S)\in\mathbb{R}^{C\times B}$, and the cross-entropy $\mathcal{L}=-\tfrac{1}{B}\sum_{i=1}^{B}\log P_{y_{i},i}$ with the label matrix $\widetilde{Y}\in[0,1]^{C\times B}$ (one-hot $Y$ or label-smoothed $\widetilde{Y}$; by default $\widetilde{Y}=Y$). The head-only gradient and proxy are
$\nabla_{W}\mathcal{L}$ $=\frac{1}{B}\,(P-\widetilde{Y})\,Z^{\top},$ (4)
$\|g\|_{F}$ $\triangleq\big\|\nabla_{W}\mathcal{L}\big\|_{F}.$
All probes detach $Z$ so that gradients do not flow into $\phi_{\psi}$.
One-shot probe (implementation).
At any training step or for a frozen checkpoint: (i) compute features $Z=\phi_{\psi}(x)$; (ii) detach $Z$; (iii) run cross-entropy through $h_{W}$ (same label-smoothing/temperature as training); (iv) zero head grads and backprop w.r.t. $W$ only (no optimizer step) to obtain $g=\nabla_{W}\mathcal{L}$; (v) record the proxy $\|g\|_{F}$. No model parameters are updated during the probe.
Scale normalization (features).
To improve cross-head comparability under feature rescaling, we also report
$\mathrm{score}_{z}\;=\;\frac{\|\nabla_{W}\mathcal{L}\|_{F}}{\|Z\|_{F}+\varepsilon},$ (5)
which is scale-aware and mitigates the effect of feature magnitude differences; we use a small stabilizer $\varepsilon$ $sss(\approx\!10^{-12})$.
Scale normalization (head).
When classifier-head magnitudes differ appreciably across checkpoints/models, we report
$\mathrm{score}_{w}\;=\;\frac{\|\nabla_{W}\mathcal{L}\|_{F}}{\|W\|_{F}+\varepsilon_{w}},$ (6)
@@
using $\varepsilon_{w}\!(\approx\!10^{-8})$ for numerical stability. Empirically, feature-scale normalization works well for Transformers/modern CNNs, while head-scale normalization is often more stable within classic CNN families (e.g., ResNet).
Detection heads (dense / query).
Let $W_{\mathrm{cls}}$ denote the per-anchor (dense) or per-query (DETR-style [5, 34, 48] classifier. Given the detector’s native assignment rule $\mathcal{A}$ (e.g., IoU-thresholding, ATSS [47], or Hungarian matching [24]), we define the classification-head probe
$\|\nabla_{W_{\mathrm{cls}}}\mathcal{L}_{\mathrm{cls}}\bigl(h_{W_{\mathrm{cls}}}(\phi_{\psi}(x)),\,\mathcal{A}(x)\bigr)\|_{F},$ (7)
where $\mathcal{L}_{\mathrm{cls}}$ is the detector’s native classification loss (e.g., CE or focal). Backbone/geometry branches are detached; unmatched queries use the standard “no-object” label.
Extending the probe to diffusion.
For a diffusion model with noise-prediction head $h_{W}(z)$ (e.g., the final $3{\times}3$ conv of a UNet), we reuse the head-only probe by backpropagating the native MSE objective on a fixed probe set:
$\mathcal{L}_{\mathrm{diff}}\;=\;\tfrac{1}{B}\!\sum_{i=1}^{B}\bigl\|\hat{\varepsilon}(x_{t_{i}},t_{i})-\varepsilon_{i}\bigr\|_{2}^{2},$ (8)
$\mathrm{score}_{w}\;=\;\frac{\bigl\|\nabla_{W}\mathcal{L}_{\mathrm{diff}}\bigr\|_{F}}{\|W\|_{F}+\varepsilon_{w}}\,$, (9)
We detach all non-head branches (or zero out their graidents) and record the head-only gradient norm per evaluation step. In practice, we evaluate on a fixed set of B=256B\!=\!256 images with timesteps $t\!\sim\!\mathrm{Uniform}([0.1T,0.7T])$, repeat $K$ times and aggregate by median, and optionally apply an EMA (exponential moving average) [4, 36] to stabilize the trajectory. No parameters are updated during the probe [18]. Unless stated otherwise, sampling for evaluation uses DDIM [40] with $\eta{=}0$ under EMA weights (decay $0.999$) for efficiency and stability.
Validation-free selection rule.
@@
Given checkpoints $\{\theta_{t}\}$ from one run, select $\hat{t}=\arg\min_{t}m(\theta_{t})$, where $m\in\{\|g\|_{F},\,\mathrm{score}_{z},\,\mathrm{score}_{w}\}$. For reporting we compare Top-1/mAP at $\hat{t}$ with the oracle best validation checkpoint; the selection itself never uses validation labels.
@@
The probe backpropagates only through the head. For $W\!\in\!\mathbb{R}^{C\times d}$ and batch size $B$, the extra FLOPs are $\mathcal{O}(CdB)$ and the dominant extra memory is the head gradient ($C\times d$) plus logits/probabilities ($C\times B$), negligible relative to a full validation pass.
@@ 5.5 Implementation Details
Unless otherwise specified we use PyTorch’s reduction=’mean’, so per-batch gradients are averaged over $B$ and the head-gradient scale is largely insensitive to batch size. Throughout, the probe backpropagates only w.r.t. $W$ (the classifier head), consistent with Sec. 4.
6 Results
@@ 6.3 Instance Segmentation
7 Application: Validation-Free Checkpoint Selection
Training deep neural networks typically involves saving multiple checkpoints and selecting the best one based on validation performance. This process requires repeated evaluation on held-out data, which can be expensive for large-scale deployments. We demonstrate that our gradient norm metric enables effective checkpoint selection without validation data. Our selection rule never uses validation labels at selection time; for reporting we still evaluate Top-1/mAP and report correlations on standard validation splits.
@@ 7.1 Universal Configuration Baseline
We select checkpoints by minimizing the EMA-smoothed head-gradient score (Section 4) over a tail window; unless noted we use k=3k\!=\!3 and s=80s\!=\!80 last steps. For reporting, we compare the selected checkpoint against the tail oracle and define the selection gap $\Delta\!=\!\mathrm{Acc}_{\text{oracle}}-\mathrm{Acc}_{\text{selected}}$ ($0\%$ indicates a perfect selection). Under this universal configuration, the average gap across seven architectures is $4.24\%\!\pm\!2.00\%$. Lightweight per-architecture tuning reduces the mean gap to $1.12\%$ (Fig. 7). Sensitivity to the EMA span $k$ and tail-window size $s$ (grid over $\{1,3,5,9\}\times\{60,80,100\}$) is reported in supplementary material.
@@
8 Application: Validation-free Family-wise Pre-screening (ResNet)
Setup. We measure a single-batch, head-only gradient on ImageNet-1k (training split; features detached), and rank checkpoints/models without using validation data at selection time. Within the ResNet family (18/34/50/101/152), we report the correlation between the head-gradient proxy and validation Top-1, following equation 6.
@@ 8 Application: Validation-free Family-wise Pre-screening (ResNet)
9 Application: Head-gradient Tracks Progress and Checkpoint Selection (Diffusion)
@@ 10 Discussion and Limitations
Why does it work? When features are more linearly separable for the task at hand, the classifier head operates in a flatter local landscape and receives smaller training signals, yielding smaller head gradients; empirically this aligns with better generalization.
Scope and boundary conditions. Our probe is designed for settings with an explicit classification head and conventional training hyperparameters. Extremely small/large batches, aggressive or unusual regularization, or tasks without a clear classification head may alter the gradient geometry; in such cases, simple adaptations (e.g., modest batch sizing, temperature/scale normalization, or including a lightweight localization term) can be used.
Labels and data usage. The method is supervised but label usage is minimal (one batch) and does not require any validation data.
Relation to NAS and scope. On heterogeneous, cross-family pools with small $n$, our single-batch head-gradient signal does not reliably rank candidates (e.g., Spearman $\rho=-0.08$ with wide 95% CIs overlapping zero), and Recall@K approaches random. This suggests that the probe is better suited to within-family comparisons and within-run checkpoint selection, where we observe strong and stable correlations. We therefore view the probe as a low-cost component for candidate pre-screening and early pruning inside NAS pipelines, rather than a stand-alone cross-family NAS method. A full end-to-end NAS benchmark is beyond our scope and left to future work; see supplementary material for diagnostics and extended results.

### v2

# Model-Centric Diagnostics: A Framework for Internal State Readouts

@@ Abstract
We present a model-centric diagnostic framework that treats training state as a latent variable and unifies a family of internal readouts—head-gradient norms, confidence, entropy, margin, and related signals—as anchor-relative projections of that state. A preliminary version of this work introduced a head-gradient probe for checkpoint selection. In this version, we focus on the unifying perspective and structural diagnostics; full algorithmic details, theoretical analysis, and experimental validation will appear in a forthcoming paper.
We outline the conceptual scaffold: any prediction head induces a local loss landscape whose geometry (gradient magnitude, curvature, sharpness) reflects how well the upstream features are aligned with the task. Different readout choices—gradient norms, softmax entropy, predictive margin—correspond to different projections of this geometry, each with complementary strengths. The framework suggests that checkpoint selection, early stopping, and lightweight architecture pre-screening can all be viewed as querying the same underlying state through different lenses.
Illustrative experiments on ImageNet classification and COCO detection/segmentation hint at the practical potential; rigorous benchmarks and ablations are deferred to the full paper.
11footnotetext: Disclaimer: This manuscript presents a conceptual framework and differs substantially from a forthcoming submission that provides full algorithmic details, theoretical analysis, and comprehensive experimental validation. The preliminary results shown here are illustrative only.
@@ 1 Introduction
Modern training loops heavily lean on held-out validation sets to decide when to stop and which checkpoint or model to keep. This practice costs extra compute for periodic evaluation, can be impossible when labels are scarce or private, and may be unreliable under distribution shift. We revisit a minimalistic question: can one predict model quality using internal state readouts—signals derived from the model itself—without any validation data at selection time?
A unifying perspective.
We argue that seemingly disparate diagnostics—gradient norms, softmax confidence, entropy, margin, and curvature-based measures—can all be understood as anchor-relative projections of a latent training state. When features become more linearly separable, the classifier head operates in a flatter local landscape. This geometry manifests through multiple observable proxies: smaller head gradients, higher confidence, lower entropy, and wider margins. Each readout offers a different “lens” into the same underlying state.
Framework contributions.
This paper sketches a model-centric diagnostic framework with the following elements:
  • Training state as latent variable. We treat the quality of learned representations as a hidden variable that causally influences multiple observable readouts.
  • Anchor-relative projections. Each diagnostic (gradient norm, confidence, etc.) projects the latent state onto a scalar or low-dimensional summary, analogous to different projections of a high-dimensional object.
  • Task-agnostic structure. The same conceptual scaffold applies to classification, detection, segmentation, and generative modeling, though the optimal readout may differ by task.
Scope of this version.
A preliminary version of this work focused on one specific instantiation: the head-gradient norm as a checkpoint selection signal. Here, we step back to emphasize the unifying framework rather than any single method. We include illustrative experiments to motivate the perspective, but defer rigorous benchmarks, ablations, and theoretical analysis to a forthcoming full paper.
Why this matters.
If internal readouts reliably track model quality, practitioners gain several benefits: validation-free early stopping, lightweight pre-screening of architectures, and monitoring in settings where labels are unavailable or expensive. The framework perspective suggests that these applications share a common foundation and that progress on one readout informs the others.
@@ 2 Related Work
Connections between loss landscape geometry and generalization have been studied via flatness/sharpness and Hessian spectra (e.g., 19, 18, 33, 12). Large gradients often co-occur with sharper regions of the loss; our findings align with this picture: models/checkpoints with smaller head gradients tend to exhibit better Top-1 and lower loss, suggesting improved separability of the learned features under the current head.
@@
3 Framework: Internal State Readouts
This section describes the conceptual framework that unifies various internal diagnostics as projections of training state. We present one concrete instantiation—the head-gradient norm—as an illustrative example; a complete treatment with theoretical analysis and systematic comparisons will appear in the full paper.
3.1 Training State as Latent Variable
We posit that the “quality” of learned representations at any checkpoint can be modeled as a latent variable $\mathcal{S}$ that causally influences multiple observables. As training progresses and features become more linearly separable, this latent state improves, and various readouts reflect that improvement:
  • Gradient-based: Head-gradient norms, full-model gradient norms, Fisher information traces
  • Output-based: Softmax confidence, predictive entropy, margin between top-2 classes
  • Geometry-based: Loss curvature, sharpness measures, Hessian eigenvalues
Each readout is a different projection of $\mathcal{S}$—analogous to viewing a 3D object from different angles.
3.2 Why Gradient Norms?
The gradient norm at a checkpoint reflects the local loss curvature [19, 18]. For a model with parameters $\theta$ at checkpoint $t$:
@@ 3.2 Why Gradient Norms?
The connection to generalization is well-established: flat minima (low sharpness) generalize better than sharp minima. The head-gradient norm, in particular, measures how “settled” the final classification layer is given the upstream features—a direct proxy for feature quality.
3.3 Example Instantiation: Head-Gradient Probe
As one concrete instance of the framework, we describe a head-gradient probe that has shown promise in preliminary experiments. This is presented as an illustrative example; the full paper will provide systematic comparisons across multiple readout types.
Notation.
@@ 3.3 Example Instantiation: Head-Gradient Probe
Let $f_{\theta}(x)=h_{W}(\phi_{\psi}(x))$, where $\phi_{\psi}:\mathcal{X}\!\to\!\mathbb{R}^{d}$ is a feature extractor and $h_{W}(z)=Wz$ is a linear head with $W\in\mathbb{R}^{C\times d}$. Given a mini-batch $\{(x_{i},y_{i})\}_{i=1}^{B}$, define the detached feature matrix $Z=[\phi_{\psi}(x_{1}),\ldots,\phi_{\psi}(x_{B})]\in\mathbb{R}^{d\times B}$, the probabilities $P=\mathrm{softmax}(WZ)$, and the cross-entropy $\mathcal{L}=-\tfrac{1}{B}\sum_{i=1}^{B}\log P_{y_{i},i}$. The head-only gradient norm is:
$\|g\|_{F}\triangleq\big\|\nabla_{W}\mathcal{L}\big\|_{F}=\Big\|\frac{1}{B}\,(P-Y)\,Z^{\top}\Big\|_{F}.$ (2)
Features are detached so that gradients do not flow into the backbone $\phi_{\psi}$.
Scale normalization variants.
To improve comparability across architectures, one can normalize by feature or head magnitude:
$\mathrm{score}_{z}=\frac{\|g\|_{F}}{\|Z\|_{F}+\varepsilon},\quad\mathrm{score}_{w}=\frac{\|g\|_{F}}{\|W\|_{F}+\varepsilon_{w}}.$ (3)
@@
Preliminary results suggest feature-scale normalization works well for Transformers/modern CNNs, while head-scale normalization is often more stable within classic CNN families.
Selection rule.
@@
Given checkpoints $\{\theta_{t}\}$ from one run, a simple validation-free rule is: $\hat{t}=\arg\min_{t}\|g\|_{F}(\theta_{t})$. The selection itself never uses validation labels.
@@
The probe backpropagates only through the head: $\mathcal{O}(CdB)$ FLOPs for $W\in\mathbb{R}^{C\times d}$ and batch size BB—negligible relative to a full validation pass.
@@ 4.5 Implementation Details
Unless otherwise specified we use PyTorch’s reduction=’mean’, so per-batch gradients are averaged over $B$ and the head-gradient scale is largely insensitive to batch size. Throughout, the probe backpropagates only w.r.t. $W$ (the classifier head), consistent with Sec. 3.3.
5 Illustrative Experiments
The following experiments are illustrative and motivate the framework perspective. Rigorous benchmarks with full ablations and statistical controls will appear in the forthcoming complete paper.
@@ 5.3 Instance Segmentation
6 Illustrative Use Case: Checkpoint Selection
This section illustrates how the head-gradient readout might be applied for checkpoint selection. Full validation of these claims requires the rigorous experiments planned for the complete paper.
Training deep neural networks typically involves saving multiple checkpoints and selecting the best one based on validation performance. This process requires repeated evaluation on held-out data, which can be expensive for large-scale deployments. Preliminary results suggest that gradient-based readouts may enable effective checkpoint selection without validation data. Our selection rule never uses validation labels at selection time; for reporting we still evaluate Top-1/mAP and report correlations on standard validation splits.
@@ 6.1 Universal Configuration Baseline
We select checkpoints by minimizing the EMA-smoothed head-gradient score (Section 3.3) over a tail window; unless noted we use k=3k\!=\!3 and s=80s\!=\!80 last steps. For reporting, we compare the selected checkpoint against the tail oracle and define the selection gap $\Delta\!=\!\mathrm{Acc}_{\text{oracle}}-\mathrm{Acc}_{\text{selected}}$ ($0\%$ indicates a perfect selection). Under this universal configuration, the average gap across seven architectures is $4.24\%\!\pm\!2.00\%$. Lightweight per-architecture tuning reduces the mean gap to $1.12\%$ (Fig. 7). Sensitivity to the EMA span $k$ and tail-window size $s$ (grid over $\{1,3,5,9\}\times\{60,80,100\}$) is reported in supplementary material.
@@
7 Illustrative Use Case: Family-wise Pre-screening
Setup. We measure a single-batch, head-only gradient on ImageNet-1k (training split; features detached), and rank checkpoints/models without using validation data at selection time. Within the ResNet family (18/34/50/101/152), we report the correlation between the head-gradient proxy and validation Top-1, using the weight-normalized score $\mathrm{score}_{w}$ (Sec. 3.3).
@@ 7 Illustrative Use Case: Family-wise Pre-screening
8 Illustrative Use Case: Diffusion Model Monitoring
@@ 9 Discussion and Limitations
Framework perspective. The preliminary results above are consistent with our hypothesis that multiple internal readouts—gradient norms, confidence, entropy, margin—are correlated projections of a common latent training state. When features become more linearly separable, the head gradient shrinks, confidence rises, and entropy falls. The full paper will systematically test this hypothesis with controlled comparisons across readout types.
Scope of preliminary results. The experiments shown here are illustrative and not exhaustive. We have not yet performed the rigorous ablations, hyperparameter sweeps, or statistical tests needed to make strong claims. In particular:
  • The probe is designed for settings with an explicit classification head; tasks without such structure may require adaptation.
  • Cross-family ranking (e.g., ResNet vs. ViT) remains unreliable; the probe is better suited to within-family comparisons.
  • Sensitivity to batch size, regularization, and training dynamics requires further study.
What the full paper will provide. The forthcoming complete paper will include: (1) theoretical analysis connecting gradient norms to generalization bounds; (2) systematic comparisons across multiple readout types (gradient, confidence, entropy, margin); (3) comprehensive benchmarks with proper statistical controls; (4) ablations on hyperparameters and boundary conditions; and (5) practical guidelines for deployment.